Models and Model Building Question: We have acquired new furniture for the office. The invoice for $6,000 offers two ways to pay: we can pay the entire amount by September 1, or we can pay $3,060 by September 1 and $3,000 by January 1. How does our decision depend on the interest rate at which we can invest our funds? SOLVE using excel please
Answer #1
We have acquired new furniture for the office. The invoice for $6,000 offers two ways to pay: we can pay the entire amount by September 1, or we can pay $3,060 by September 1 and $3,000 by January

P  Present value
n  number of periods
r  rate of interest
FV Future value

P    = 6000-3060 
     = 2940
Fv   = P+PNR/ 100
3000 = 2940+2940*4/12*R/100
60   = 2940*4/12*R/100
60*12/4  = 2940*R/100
180      = 2940*R/100
180/2940 = R/100
0.061224 = R/100

R = 6.1224 

Do not accept if your outside loan Interest rate < 6.12%